Two resistors, 47 ohms and 35 ohms, are connected in parallel. The current through the 47 ohm resistor is 2.0 A.  What is the total power supplied to the two resistors?Question 21 (3 points) Two resistors, 47 ohms and 35 ohms, are connected in parallel. The current through the 47 ohm resistor
Answer #1

Answer:- Total power supplied is 440 W

Explanation:-

Given:

  • Reaistor 1 : R_{1} = 47 \ \Omega
  • Reaistor 2 : R_{2} = 35 \ \Omega
  • Current through reaistor 1 : I_{1} = 2 \ A

The voltage across the resistor 1 is given by,

V_{1} = R_{1} I_{1}

V_{1} = 47 \times 2

V_{1} = 94 \ V

Since these two resistors are connected in parallel they must have same voltage across them.

V_{1} = V_{2}

Therefore the voltage across the resistor 2 will be,

V_{1} = V_{2} = 94 \ V

Thus, the current through the resistor 2 is given by,

V_{2} = R_{2} I_{2}

I_{2} = \frac{V_{2} }{R_{2} }

I_{2} = \frac{ 94 }{35}

I_{2} = 2.685 \ A

Now, the power supplied is given by,

P = I ^{2} R

The total power is sum of the power supplied to two resistors,

P = P_{1} + P_{2}

P =( I _{1} )^{2} R_{1} + ( I_{2}) ^{2} R_{2}

P =( 2 )^{2} \times 47 + ( 2.685 ) ^{2} \times 35

P = 188 + 252.3

P = 440.3

\boldsymbol{P = 440 \ W}

This is the total power supplied to the two resistors.